OpenStudy (anonymous):

convert to rectangular form r=20/(1-cos(theta))

ganeshie8 (ganeshie8):

\(\large \mathbb{r = \frac{20}{1-\cos \theta}}\)

ganeshie8 (ganeshie8):

\(\large \mathbb{r(1-\cos \theta ) = 20}\) \(\large \mathbb{r-r\cos \theta = 20}\) \(\large \mathbb{r = 20 +r\cos \theta }\)

ganeshie8 (ganeshie8):

substitute : \(r\cos \theta = x\) \(r^2 = x^2+y^2\)

ganeshie8 (ganeshie8):

\(\large \mathbb{x^2+y^2 = (20 +x)^2 }\) simplify

OpenStudy (anonymous):

thank you Ganeshie8

ganeshie8 (ganeshie8):

np :) simplify it... x^2 cancels out... its just a parabola..